# Power Battery Detection

### v1

@@ II-C Enhancing Weak Features in Complex Vision Tasks
In complex vision tasks, key targets often exhibit weak, sparse, or ambiguous features due to small object scales, low contrast, occlusion, vague boundaries, or background clutter. Such weak signals pose fundamental challenges in remote sensing, medical imaging and industrial inspection. To address this, recent studies enhance weak features via two main strategies: multi-modal and multi-structural clues fusion, both injecting complementary or contextual information into feature learning. Multi-modal fusion improves feature expressiveness by introducing multiple modalities. For example, Liu et al. [82] employ RGB-D for camouflaged object segmentation, Kim et al. [83] utilize RGB-T for transparent object segmentation, and Pang et al. [84] verify the complementarity between RGB-D and RGB-T modalities. Beyond visual signals, text serves as a strong auxiliary modality in semantically ambiguous scenes. CLIP [85] enable open-vocabulary recognition via large-scale image-text pretraining, while Pang et al. [23] use textual descriptions to generalize camouflaged object segmentation to novel categories. In the medical domain, cross-modality fusion (e.g., MRI, CT, PET) provides complementary anatomical and functional context for localizing subtle lesions [86]. Beyond modalities, structural priors such as edges, contours, and symmetry help delineate fuzzy or ambiguous targets. Edge-aware modules [87, 88, 89] guide fine-grained segmentation, while higher-level topological clues (e.g., skeletons, geometric primitives) benefit tasks like polyp segmentation [90], shadow removal [91], and transparent object detection [92]. These structures isolate meaningful patterns amid visual noise. Our work aligns with this direction by constructing a multi-clue segmentation framework that jointly leverages point, line, and number clues to enhance weak feature perception, enabling robust object discovery in X-ray battery images with complex interference.
III PBD5K Dataset
III-A Image Collection and Annotation
@@
Automated Screening. The pipeline begins with the automated filtering of raw data. This stage includes battery integrity screening and duplicate image filtering. The former employs a lightweight object detection model to discard invalid samples with incomplete views (e.g., batteries cut off at edges), overexposed images, and blank captures. The latter eliminates near-duplicate images by computing high-level visual similarity using pre-trained convolutional neural network (CNN) features. Specifically, each image is first encoded into a compact 512-dimensional representation using global average pooling over intermediate VGG [93] feature maps. To handle large-scale retrieval, we apply FAISS [94], a fast approximate nearest neighbor search library. Image pairs with Euclidean distance below a threshold are deemed duplicates. We then perform connected-component clustering on the similarity graph and retain one representative per cluster. This multi-stage screening ensures the final dataset is both structurally complete and semantically diverse.
Figure 3: The overall intelligent annotation pipeline with automated screening, annotation engine with active Learning, and multi-expert quality control. $\color[rgb]{202,0,0}{\text{Red arrows}(\longrightarrow)}$ indicate poor annotation quality, while $\color[rgb]{0,136,51}{\text{green arrows}(\longrightarrow)}$ indicate high-quality annotations that are included in the Gold Standard Set.
@@
During the entire process, human expert teams work in parallel rather than sequentially, ensuring high throughput. This active learning-driven annotation loop iteratively improves MDCNeXt and continues until convergence. Through this pipeline, we construct the PBD5K dataset, a gold-standard benchmark composed of 5,000 expertly annotated X-ray battery images, validated through automated and human-in-the-loop quality checks. As shown in Tab. II, our approach significantly improves annotation efficiency and reliability compared to traditional manual schemes.
TABLE II: Efficiency comparison between manual and intelligent annotation schemes.
@@
III-B Dataset Analysis
Hierarchical Distribution of Categories and Attributes. As shown in Fig. 4, the collected images have 2 categories, 3 shots, 9 attributes and their proportions under each shot type are statistically reported in detail. Different battery types, manufacturing and packaging processes produce different thicknesses of batteries. Even if the DR device is calibrated, it still has different degrees of penetration for different thicknesses, thereby resulting in Clear and Blur X-ray images. The imaging field-of-view include close shot (CS), medium shot (MS) and long shot (LS). Each shot type further contains samples labeled with multiple common attributes: pure plate (P), tilted plate (T), aberrant plate (A), illumination interference (II), plate interference (PI), bifurcation interference (BI), tray interference (TRI), tab interference (TAI), and separator interference (SI). The quantitative proportions are marked beside each attribute. We can see that P only appears in the Clear category, II is exclusively found in the CS view, while SI does not occur in the CS view. The detailed definitions of these attributes are given in Tab. I, and visual samples are presented in Fig. 1.
@@ III-B Dataset Analysis
III-C Evaluation Metrics
According to the number and overhang criteria adopted by the manufacturers, we design eight complementary metrics to quantitatively evaluate the performance of algorithms. Specifically, they are number mean absolute error (AN-MAE, CN-MAE), number accuracy (AN-ACC, CN-ACC, PN-ACC) and position mean absolute error (AL-MAE, CL-MAE and OH-MAE) for the anode level, cathode level, and pair level, respectively. Let $n_{i}^{x}$ and $\hat{n}_{i}^{x}$ represent the predicted and ground truth plate numbers for the $i$-th image at level $x\in\{\text{anode, cathode, pair}\}$, and $p_{i,j}^{y}$, $\hat{p}_{i,j}^{y}$ represent the coordinates of the $j$-th plate at level $y\in\{\text{anode, cathode}\}$ for image $i$ with resolution $H_{i}\times W_{i}$. The eight evaluation metrics are defined as follows:
@@ III-C Evaluation Metrics
where $N$ is the total number of test samples and $N_{p}$ is the number of samples with correctly predicted plate numbers. Note that, for an image, only when its number of AN or CN is predicted with 100% accuracy, we can further evaluate the corresponding position MAE. Coordinates must be sorted before alculating AL-MAE, CL-MAE and OH-MAE.
@@
IV Proposed Framework
IV-A Overview
As shown in Fig. 7, the MDCNeXt adopts the encoder-decoder structure with a ResNet-50 [95] backbone. The encoder initially extracts five-level features, and the prompt and current pipelines share the same weights. The prompt-filtered state space module (PFSSM) uses the prompt features to filter out interfering information from the current features and generate pure plate appearance features. The point predictor outputs the coarse point segmentation map after passing through five decoder blocks. The line predictor and counting predictor assist in point prediction at the low-level detail and semantic levels, respectively. Finally, the density-aware reordering state space module (DRSSM) carefully refines the coarse point segmentation map and features to obtain a refined point map prediction.
IV-B Prompt-filtered State Space Module
@@
IV-C Multi-Dimensional Decoder
Our multi-dimensional decoder consists of a point predictor, counting predictor, and line predictor. The point predictor adopts a general U-shaped architecture [11, 10], progressively fusing features output by each PFSSM to generate a coarse point map highly responsive to point locations. These coarse anode and cathode maps are subsequently used by both the counting and line predictors. As shown in Fig. 9, it illustrates counting predictor and line predictor.
@@ IV-C Multi-Dimensional Decoder
where US(·) is the upsampling operation. Next, $F_{current}^{e1,2}$ is separately computed with $M^{anode}_{p}$ and $M^{cathode}_{p}$ in the form of residual [95] and predict line segmentation maps:
@@
IV-D Density-aware Reordering State Space Module
@@
IV-E Label Generation and Supervision
For point segmentation, a direct strategy of generating the ground truth of point mask is to use circular regions of the same radius for each plate. As shown in the $1^{st}$ - $3^{rd}$ columns of Fig. 11, the point masks are yielded under the radius of $1$, 33, and $5$ pixels, respectively. It can be seen that the fixed radius often produces very dense or sparse point masks for different types of batteries, which increases the difficulty of point segmentation. To this end, we design a novel distance-adaptive label generation strategy. It computes the diameter of each point according to the distance of adjacent plates. As shown in the $4^{th}$ - $6^{th}$ columns of Fig. 11, the point masks are generated under the diameter of $0.1x$, $0.3x$, and $0.5x$ distance, respectively. The labels of counting and line segmentation can be separately obtained by the connected component analysis and connecting each endpoint in the point mask in sequence. For the point and line segmentation predictors, we use the weighted IoU loss and binary cross entropy (BCE) loss, which have been widely adopted in segmentation task. We use the same definitions as in [96, 97, 98, 56]. For the counting predictor, we adopt the common L1 loss. The total training loss is written as follows:
@@ IV-E Label Generation and Supervision
V Experiments
V-A Implementation Details
@@
TABLE III: Quantitative comparison of different methods. $\uparrow$ and $\downarrow$ indicate that the larger scores and the smaller ones are better, respectively. The best scores are highlighted in red. “—” represents that the results are not available because these methods can not provide coordinate information or their prediction accuracy of the number of plates is zero. “Average” refers to the average scores across the Regular, Difficult, and Tough test splits.
@@
 [29]  [30]  [31]  [53]  [43]  [49]  [35]  [38]  [36]  [39]  [41] CVPR 2024 [21] (Ours)
@@
V-B Comparison with Corner Detection, Crowd Counting, General and Tiny Object Detection Solutions
@@
Quantitative Comparison. As shown in Tab. III, we conduct a comprehensive comparison across various representative methods that serve as potential solutions to the PBD task. The evaluation spans regular, difficult, and tough splits, and we further report the average scores across them. The key observations are as follows: I) Our segmentation-based MDCNeXt and MDCNet achieve state-of-the-art performance across different splits under all metrics. It confirms that modeling the PBD task as a visual point segmentation problem is a correct solution. Moreover, MDCNeXt further outperforms MDCNet with average gains exceeding 18%, 17%, and 20% on AN-ACC, CN-ACC, and PN-ACC, respectively. II) The complementarity among different evaluation metrics highlights the necessity of using a multi-metric evaluation scheme for PBD. For example, tiny object detection methods have AN-MAE and CN-MAE scores close to ours (within 2 units) on the regular split. However, their AN-ACC and CN-ACC drop by more than 20% compared to MDCNeXt. Besides, the OH-MAE metric may fail to capture systematic shifts. When both AN and CN position predictions are slightly displaced in the same direction, OH-MAE might still appear low. In the tough split, other methods show AL-MAE and CL-MAE scores that are more than twice their OH-MAE, exposing serious localization issues. Therefore, combining multiple metrics gives a more reliable and precise evaluation. This is consistent with the strict spatial accuracy requirements of the PBD task, where even small errors are magnified and “almost correct” results are not acceptable. III) MDCNeXt shows strong robustness and stability across all splits. Its AL-/CL-MAE values fluctuate within only 0.3. In contrast, other object detection methods suffer from significant accuracy degradation as the number of plates sheets increases and the accumulation of interference types. In the tough split, their AL-MAE and CL-MAE scores exceed 15, reflecting a strong decline in positional precision compared to the regular split.

### v2

@@ 2.3 Enhancing Weak Features in Complex Vision Tasks
In complex vision tasks, key targets often exhibit weak, sparse, or ambiguous features due to small object scales, low contrast, occlusion, vague boundaries, or background clutter. Such weak signals pose fundamental challenges in remote sensing, medical imaging and industrial inspection. To address this, recent studies enhance weak features via two main strategies: multi-modal and multi-structural clues fusion, both injecting complementary or contextual information into feature learning. Multi-modal fusion improves feature expressiveness by introducing multiple modalities. For example, Liu et al. RGBDCOD employ RGB-D for camouflaged object segmentation, Kim et al. Transpose utilize RGB-T for transparent object segmentation, and Pang et al. CAVER_RGBDSOD verify the complementarity between RGB-D and RGB-T modalities. Beyond visual signals, text serves as a strong auxiliary modality in semantically ambiguous scenes. CLIP CLIP enable open-vocabulary recognition via large-scale image-text pretraining, while Pang et al. OVCOS use textual descriptions to generalize camouflaged object segmentation to novel categories. In the medical domain, cross-modality fusion (e.g., MRI, CT, PET) provides complementary anatomical and functional context for localizing subtle lesions Multi-modality_fusion_survey . Beyond modalities, structural priors such as edges, contours, and symmetry help delineate fuzzy or ambiguous targets. Edge-aware modules Inf-Net ; CFANet_Polyp ; EGNet guide fine-grained segmentation, while higher-level topological clues (e.g., skeletons, geometric primitives) benefit tasks like polyp segmentation GSNet_Polyp , shadow removal DAS_shadow_removal , and transparent object detection Trans2Seg_Transparent . These structures isolate meaningful patterns amid visual noise. Our work aligns with this direction by constructing a multi-clue segmentation framework that jointly leverages point, line, and number clues to enhance weak feature perception, enabling robust object discovery in X-ray battery images with complex interference.
3 PBD5K Dataset
3.1 Image Collection and Annotation
@@
Automated Screening. The pipeline begins with the automated filtering of raw data. This stage includes battery integrity screening and duplicate image filtering. The former employs a lightweight object detection model to discard invalid samples with incomplete views (e.g., batteries cut off at edges), overexposed images, and blank captures. The latter eliminates near-duplicate images by computing high-level visual similarity using pre-trained convolutional neural network (CNN) features. Specifically, each image is first encoded into a compact 512-dimensional representation using global average pooling over intermediate VGG VGG feature maps. To handle large-scale retrieval, we apply FAISS FAISS , a fast approximate nearest neighbor search library. Image pairs with Euclidean distance below a threshold are deemed duplicates. We then perform connected-component clustering on the similarity graph and retain one representative per cluster. This multi-stage screening ensures the final dataset is both structurally complete and semantically diverse.
Figure 4: The overall intelligent annotation pipeline with automated screening, annotation engine with active Learning, and multi-expert quality control. Red arrows ($\longrightarrow$) indicate poor annotation quality, while green arrows ($\longrightarrow$) indicate high-quality annotations that are included in the Gold Standard Set.
@@
During the entire process, human expert teams work in parallel rather than sequentially, ensuring high throughput. This active learning-driven annotation loop iteratively improves MDCNeXt and continues until convergence. Through this pipeline, we construct the PBD5K dataset, a gold-standard benchmark composed of 5,000 expertly annotated X-ray battery images, validated
through automated and human-in-the-loop quality checks. As shown in Tab. 2, our approach significantly improves annotation efficiency and reliability compared to traditional manual schemes.
Table 2: Efficiency comparison between manual and intelligent annotation schemes.
@@
3.2 Dataset Analysis
Hierarchical Distribution of Categories and Attributes. As shown in Fig. 5, the collected images have 2 categories, 3 shots, 9 attributes and their proportions under each shot type are statistically reported in detail. Different battery types, manufacturing and packaging processes produce different thicknesses of batteries. Even if the DR device is calibrated, it still has different degrees of penetration for different thicknesses, thereby resulting in Clear and Blur X-ray images. The imaging field-of-view include close shot (CS), medium shot (MS) and long shot (LS). Each shot type further contains samples labeled with multiple common attributes: pure plate (P), tilted plate (T), aberrant plate (A), illumination interference (II), plate interference (PI), bifurcation interference (BI), tray interference (TRI), tab interference (TAI), and separator interference (SI). The quantitative proportions are marked beside each attribute. We can see that P only appears in the Clear category, II is exclusively found in the CS view, while SI does not occur in the CS view. The detailed definitions of these attributes are given in Tab. 1, and visual samples are presented in Fig. 2.
@@ 3.2 Dataset Analysis
3.3 Evaluation Metrics
According to the number and overhang criteria adopted by the manufacturers, we design eight complementary metrics to quantitatively evaluate the performance of algorithms. Specifically, they are number mean absolute error (AN-MAE, CN-MAE), number accuracy (AN-ACC, CN-ACC, PN-ACC) and position mean absolute error (AL-MAE, CL-MAE and OH-MAE) for the anode level, cathode level, and pair level, respectively. Let $n_{i}^{x}$ and $\hat{n}_{i}^{x}$ represent the predicted and ground truth plate numbers for the $i$-th image at level x{anode, cathode,
pair
[…]
, and $p_{i,j}^{y}$, $\hat{p}_{i,j}^{y}$ represent the coordinates of the $j$-th plate at level $y\in\{\text{anode, cathode}\}$ for image $i$ with resolution $H_{i}\times W_{i}$. The eight evaluation metrics are defined as follows:
@@ 3.3 Evaluation Metrics
where $N$ is the total number of test samples and $N_{p}$ is the number of samples with correctly predicted plate numbers. Note that, for an image, only when its number of AN or CN is predicted with 100% accuracy, we can further evaluate the corresponding position MAE. Coordinates must be sorted before calculating AL-MAE, CL-MAE and OH-MAE.
@@
4 Proposed Framework
4.1 Overview
As shown in Fig. 8, the MDCNeXt adopts the encoder-decoder structure with a ResNet-50 Resnet backbone. The encoder initially extracts five-level features, and the prompt and current pipelines share the same weights. The prompt-filtered state space module (PFSSM) uses the prompt features to filter out interfering information from the current features and generate pure plate appearance features. The point predictor outputs the coarse point segmentation map after passing through five decoder blocks. The line predictor and counting predictor assist in point prediction at the low-level detail and semantic levels, respectively. Finally, the density-aware reordering state space module (DRSSM) carefully refines the coarse point segmentation map and features to obtain a refined point map prediction.
4.2 Prompt-filtered State Space Module
@@
4.3 Multi-Dimensional Decoder
Our multi-dimensional decoder consists of a point predictor, counting predictor, and line predictor. The point predictor adopts a general U-shaped architecture Unet ; FPN , progressively fusing features output by each PFSSM to generate a coarse point map highly responsive to point locations. These coarse anode and cathode maps are subsequently used by both the counting and line predictors. As shown in Fig. 10, it illustrates counting predictor and line predictor.
@@ 4.3 Multi-Dimensional Decoder
where US(·) is the upsampling operation. Next, $F_{current}^{e1,2}$ is separately computed with $M^{anode}_{p}$ and $M^{cathode}_{p}$ in the form of residual Resnet and predict line segmentation maps:
@@
4.4 Density-aware Reordering State Space Module
@@
4.5 Label Generation and Supervision
For point segmentation, a direct strategy of generating the ground truth of point mask is to use circular regions of the same radius for each plate. As shown in the $1^{st}$ - $3^{rd}$ columns of Fig. 12, the point masks are yielded under the radius of $1$, 33, and $5$ pixels, respectively. It can be seen that the fixed radius often produces very dense or sparse point masks for different types of batteries, which increases the difficulty of point segmentation. To this end, we design a novel distance-adaptive label generation strategy. It computes the diameter of each point according to the distance of adjacent plates. As shown in the $4^{th}$ - $6^{th}$ columns of Fig. 12, the point masks are generated under the diameter of $0.1x$, $0.3x$, and $0.5x$ distance, respectively. The labels of counting and line segmentation can be separately obtained by the connected component analysis and connecting each endpoint in the point mask in sequence. For the point and line segmentation predictors, we use the weighted IoU loss and binary cross entropy (BCE) loss, which have been widely adopted in segmentation tasks. We use the same definitions as in SPNet ; F3Net ; PraNet ; MSNet_Polyp . For the counting predictor, we adopt the common L1 loss. The total training loss is written as follows:
@@ 4.5 Label Generation and Supervision
5 Experiments
5.1 Implementation Details
@@
Table 3: Quantitative comparison of different methods. $\uparrow$ and $\downarrow$ indicate that the larger scores and the smaller ones are better, respectively. The best scores are highlighted in red. “—” represents that the results are not available because these methods can not provide coordinate information or their prediction accuracy of the number of plates is zero. “Average” refers to the average scores across the Regular, Difficult, and Tough test splits.
@@
 Harris  Shitomasi  Sub-Pixel  BL  CUT_cc  IOCFormer  RetinaNet  DETR  Yolov10  C3Det  HS-FPN CVPR 2024 MDCNet (Ours)
@@
5.2 Comparison with Corner Detection, Crowd Counting, General and Tiny Object Detection Solutions
@@
Quantitative Comparison. As shown in Tab. 3, we conduct a comprehensive comparison across various representative methods that serve as potential solutions to the PBD task. The evaluation spans regular, difficult, and tough splits, and we further report the average scores across them. The key observations are as follows: I) Our segmentation-based MDCNeXt and MDCNet achieve state-of-the-art performance across different splits under all metrics. It confirms that modeling the PBD task as a visual point segmentation problem is a correct solution. Moreover, MDCNeXt further outperforms MDCNet with average gains exceeding 18%, 17%, and 20% on AN-ACC, CN-ACC, and PN-ACC, respectively. II) The complementarity among different evaluation metrics highlights the necessity of using a multi-metric evaluation scheme for PBD. For example, tiny object detection methods have AN-MAE and CN-MAE scores close to ours (within 2 units) on the regular split. However, their AN-ACC and CN-ACC drop by more than 20% compared to MDCNeXt. Besides, the OH-MAE metric may fail to capture systematic shifts. When both AN and CN position predictions are slightly displaced in the same direction, OH-MAE might still appear low. In the tough split, other methods show AL-MAE and CL-MAE scores that are more than twice their OH-MAE, exposing serious localization issues. Therefore, combining multiple metrics gives a more reliable and precise evaluation. This is consistent with the strict spatial accuracy requirements of the PBD task, where even small errors are magnified and “almost correct” results are not acceptable. III) MDCNeXt shows strong robustness and stability across all splits. Its AL-/CL-MAE values fluctuate within only 0.3. In contrast, other object detection methods suffer from significant accuracy degradation as the number of plates sheets increases and the accumulation of interference types. In the tough split, their AL-MAE and CL-MAE scores exceed 15, reflecting a strong decline in positional precision compared to the regular split.
Figure 14: Visual comparison with other general/tiny object detection-based HS-FPN ; CFINet ; Yolov10 ; DETR ; RetinaNet , counting-based BL ; CUT_cc ; IOCFormer , corner detection-based Sub-Pixel ; Shitomasi ; Harris solutions. We directly visualize the predicted results (Ours: Segmentation map, General/Tiny object detection methods: Bounding box, Counting methods: Density map, Corner detection methods: Corner map) without any post-processing operations. Black dashed boxes ([--]) highlight the regions with obvious prediction failures. Best viewed zoomed in.
Table 4: Quantitative comparison of image segmentation methods. All ten metrics are widely used across various segmentation branches SegFormer ; ZoomNet ; PraNet ; CDCU-Spider , with details introduced in the survey GateNetv2 . $\uparrow$ and $\downarrow$ indicate that higher and lower values are better, respectively. Best scores are highlighted in red. MDCNeXt’s relative improvements over MDCNet are highlighted in green. “Average” denotes the mean scores across Regular, Difficult, and Tough test splits. For prompt-based generalist models (i.e., SAM 2 and SegGPT), we use the same pure battery images and their corresponding masks as those in MDCNeXt as prompts to generate predictions for the test set.
Regular SAM 2 SAM2 ICLR 2025 Hiera-B+ Hiera 0.9802 0.2473 0.2213 0.1979 0.5932 0.7026 0.5121 0.2722 0.3541 0.0177
SegGPT SegGPT ICCV 2023 ViT-L ViT 0.9818 0.2163 0.1958 0.1932 0.5541 0.7319 0.5551 0.1958 0.3880 0.0182
DeepLabV3+ DeepLabV3+ ECCV 2018 ResNet-101 Resnet 0.9867 0.5078 0.4278 0.4106 0.6438 0.9124 0.5920 0.4379 0.2711 0.0147
SegFormer SegFormer NeurIPS 2021 MiT-B5 SegFormer 0.9848 0.5435 0.4326 0.4020 0.6726 0.9074 0.6020 0.4541 0.2820 0.0134
Spider CDCU-Spider ICML 2024 ConvNeXt-L ConvNeXt 0.9888 0.6425 0.5970 0.6212 0.7836 0.9425 0.6820 0.6278 0.1422 0.0097
ZoomNeXt ZoomNeXt TPAMI 2024 ResNet-50 Resnet 0.9888 0.6320 0.6073 0.6320 0.7825 0.9617 0.6923 0.6342 0.1302 0.0092
MDCNet MDCNet CVPR 2024 ResNet-50 Resnet 0.9909 0.6730 0.6635 0.7083 0.7912 0.9687 0.7432 0.7023 0.1107 0.0081
MDCNeXt ResNet-50 Resnet 0.9924 0.6979 0.6971 0.7337 0.8113 0.9786 0.7857 0.7248 0.1101 0.0076
Difficult SAM 2 SAM2 ICLR 2025 Hiera-B+ Hiera 0.9836 0.1989 0.1782 0.1678 0.5029 0.7302 0.5780 0.1011 0.4474 0.0187
SegGPT SegGPT ICCV 2023 ViT-L ViT 0.9817 0.1499 0.1359 0.1337 0.5281 0.6867 0.5335 0.1347 0.4233 0.0182
DeepLabV3+ DeepLabV3+ ECCV 2018 ResNet-101 Resnet 0.9865 0.3691 0.3742 0.3542 0.4109 0.9102 0.5532 0.3801 0.3080 0.0142
SegFormer SegFormer NeurIPS 2021 MiT-B5 SegFormer 0.9888 0.3390 0.3512 0.3210 0.3902 0.8921 0.5721 0.3670 0.3303 0.0153
Spider CDCU-Spider ICML 2024 ConvNeXt-L ConvNeXt 0.9900 0.4412 0.4740 0.4412 0.6021 0.9314 0.6023 0.4921 0.2344 0.0129
ZoomNeXt ZoomNeXt TPAMI 2024 ResNet-50 Resnet 0.9888 0.5012 0.5079 0.4936 0.6785 0.9508 0.6289 0.5182 0.2161 0.0117
MDCNet MDCNet CVPR 2024 ResNet-50 Resnet 0.9908 0.5652 0.5420 0.5372 0.7084 0.9538 0.6709 0.5690 0.1761 0.0094
MDCNeXt ResNet-50 Resnet 0.9913 0.6252 0.6245 0.6611 0.7726 0.9758 0.7459 0.6522 0.1461 0.0086
Tough SAM 2 SAM2 ICLR 2025 Hiera-B+ Hiera 0.9840 0.1312 0.1102 0.0981 0.4980 0.7022 0.5320 0.0921 0.4467 0.0186
SegGPT SegGPT ICCV 2023 ViT-L ViT 0.9819 0.0914 0.0824 0.0776 0.5092 0.6022 0.5158 0.0790 0.4579 0.0180
DeepLabV3+ DeepLabV3+ ECCV 2018 ResNet-101 Resnet 0.9878 0.2712 0.1997 0.1939 0.2892 0.7920 0.3702 0.3083 0.3999 0.0178
SegFormer SegFormer NeurIPS 2021 MiT-B5 SegFormer 0.9890 0.2521 0.2001 0.2032 0.2782 0.7808 0.3840 0.3012 0.3911 0.0173
Spider CDCU-Spider ICML 2024 ConvNeXt-L ConvNeXt 0.9896 0.3831 0.3221 0.3353 0.4483 0.9312 0.5201 0.4032 0.2411 0.0146
ZoomNeXt ZoomNeXt TPAMI 2024 ResNet-50 Resnet 0.9888 0.4378 0.3792 0.3511 0.5189 0.9433 0.5811 0.4732 0.2235 0.0129
MDCNet MDCNet CVPR 2024 ResNet-50 Resnet 0.9890 0.4735 0.4012 0.3832 0.5733 0.9511 0.6021 0.5034 0.2190 0.0121
MDCNeXt ResNet-50 Resnet 0.9903 0.5641 0.5633 0.5793 0.7353 0.9745 0.7153 0.5872 0.1831 0.0097
Average SAM 2 SAM2 ICLR 2025 Hiera-B+ Hiera 0.9829 0.1884 0.1663 0.1520 0.5245 0.7136 0.5455 0.1421 0.4231 0.0184
SegGPT SegGPT ICCV 2023 ViT-L ViT 0.9818 0.1472 0.1332 0.1300 0.5284 0.6697 0.5331 0.1316 0.4259 0.0181
DeepLabV3+ DeepLabV3+ ECCV 2018 ResNet-101 Resnet 0.9870 0.3717 0.3289 0.3145 0.4297 0.8708 0.5012 0.3707 0.3296 0.0155
SegFormer SegFormer NeurIPS 2021 MiT-B5 SegFormer 0.9878 0.3622 0.3210 0.3020 0.4250 0.8584 0.5161 0.3672 0.3384 0.0155
Spider CDCU-Spider ICML 2024 ConvNeXt-L ConvNeXt 0.9896 0.4734 0.4543 0.4517 0.5968 0.9342 0.5950 0.4970 0.2129 0.0127
ZoomNeXt ZoomNeXt TPAMI 2024 ResNet-50 Resnet 0.9888 0.5134 0.4899 0.4810 0.6513 0.9511 0.6290 0.5328 0.1965 0.0115
MDCNet MDCNet CVPR 2024 ResNet-50 Resnet 0.9902 0.5619 0.5256 0.5291 0.6840 0.9567 0.6662 0.5811 0.1738 0.0100
MDCNeXt ResNet-50 Resnet 0.9912 0.6232 0.6225 0.6521 0.7699 0.9761 0.7458 0.6489 0.1494 0.0087